find the SLOPE: between (-3,5) and (-1,-3)
Answer by lenny460(1073)

Find the SLOPE: between (-3,5) and (-1,-3)

We find the slope using:

m = (y2 - y1) / (x2 - x1)
where
x1 = -3
y1 = 5
x2 = -1 and
y2 = -3

Therefore:
m = (-3 - 5) / (-1 - (-3)
-3 - 5 = -8
-1 - (-3) = -1 + 3 = 2

Therefore:
-8 / 2

The Slope :
-4



Answer:
-4







Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: obuong3@aol.com